Minutes BWS Upgrade: Electro-Mechanical Design Meeting #7 -- 6th November 2015 --

Participants: Raymond Veness, Bernd Dehning, William Andreazza, Dmitri Gudkov, Paolo Magagnin, Pierre-Jean Lapray, Jose Luis Sirvent

Agenda ( http://indico.cern.ch/event/460286/ )

PDF Version here        Bernd's notes here

1. Follow Up Open Actions:

  • Stepped vacuum chamber: Decided to keep 3 steps, new studies would be required to validate different design.
  • Feedthrough Extension: Distances checked, no problem with extended feedthrough for integration.
  • Airgap for resolver 0.25: Company contacted to check if is possible to increase.
  • Absolute positioning of scanner: Accuracy Needed by <200um Referenced to nearby quads.

2. Communications:

  • Journey to Alxion: Dmitry and Jonathan will go. (30th-Nov + 1st-Dec) The visit is still under question. Confirmation needed from CERN legal and purchasing services.
  • Metallic disk shows better performance in non-reflective areas than glass disk, but roughness must be improved.
  • SPS BWS Controller Installation
    • Firmware: Ok
    • Driver: Not yet ready
    • Testings shifted one week Week 46 --> Week 47

3. Integration considerations and actuator status. (D.Gudkov)

  • Meeting with RF Impedance Team:
    • Required Resistance of Wire on SPS - (1Kohm -10Kohm).
    • Request of CF16 for Coupler - Already implemented on the 3D design. The interface changed from CF16 to CF40
  • Meeting with Vacuum Team:
    • Additional pump for BWS in PSB required - Based on simulations with SPS proto
    • Agreed for 3D vacuum simulations with PSB design. Model needs to be provided.
    • CF DN100 interface required on each tank for vacuum pump - Already implemented on the 3D design.
    • Budget for new pump need to be foreseen ~12,5KCHF/pump
    • Materials testings on January'16
    • BE-OP required to understand better vacuum parameters on PSB.
  • Design Status:
    • Four viewports integrated on design: No problem for integration, mounting/dismounting.
    • CF16 for RF coupler on the "back part" OK for integration but need to be changed by CF40 flange.
    • Schedule:
      • Drawing Approval --> December 2015
      • Launch Production --> January 2016
  • Kollmorgen Test-Bench: Ready
  • Alxion Special motor: Some pieces sent for cleaning and sent back.
  • Comments:
    • Vacuum acceptance tests: Start with non-standard components (forks, Kapton wires, rotor, bearings...)
    • Need to re-define adjustment ranges and positions with Jonathan: To be seen on next meeting 19_11_2015.This point was discussed offline and solution was found.

4. Wire Scanner Shaft: Investigations and other news. (P.Magagnin)

  • Space Reservation:
    • Requirements of optical fibre rack and other racks to be specified.
    • Already specified details about required cabling to be pulled.
  • Aperture: Same for both locations and for both planes in PSB.
  • Shaft Optimization:
    • Inertia increase from 2.94 to 5.43e-4 Kg*m2
    • Long discussion about screwing both parts of the shaft together
      • Paolo will give an estimate of forces in the screws

5. AOB

  • How to proceed with BWS design for SPS?
    • PS, Booster & LHC same can be the same device with different forks?
    • SPS needs a different shaft lenght.
    • Currently for SPS there are pairs of scanners, requested to optimize. Scanners new placements optimized with respect to Beta function.
    • Discussion with Hannes required
  • Request about putting a BWS on the PS
    • Not yet foreseen in schedule
    • Current PSB design may work on the two planes (William).
    • Schedule for YETS17?
  • Discussion about wire mounting:
    • Procedure needs to start being considered
    • Proposal: Direct soldering the wire to fork
      • Vacuum considerations: Type of soldering element needed? Tin, Silver...
    • Tooling development for wire mounting is required (William already took it into consideration)
  • Type of wire for the PSB:
    • Mathew could take a look for proposals.
    • Multiwire seems better from the heat point of view.
  • SPS Optical Calibration Test-Bench:
    • Modification required for table usage with laboratory proto.
  • BWS Laboratory Prototype:
    • Plexiglass cover needed for security.
    • Optical fibre support needed

6. New Actions:

Action Responsible
Provide material for Vacuum Tests Dmitry, William
Contact resolver company to check if possible bigger airgap Jonathan
Inertia update to be prepared Paolo
Forces on shaft screws to be calculated
Run tests on new disk samples (done) Jose
Validate new lenses and PSB optical position sensor distances (done)
Check new disk holder on BWS prototype
Contact Hannes concerning aperture for the SPS scanners (done) Bernd
Brake and angular adjustments (to be discussed on next meeting)
Carbon Wire outside orders EDH
Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2015-11-24 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    BWSUpgrade All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback